# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ III.3 Quantum Hall Family of Phonons
Figure 20: Various topological phonon states in 2D honeycomb lattices. (a) A honeycomb lattice hosts a 2D topological Dirac phonon phase, featuring a pair of Dirac cones at the Brillouin zone corners $K$ and $K^{\prime}$. (b)-(d) By breaking different symmetries, the Dirac cones can be gapped, leading to the emergence of topologically distinct phases, including (b) quantum (anomalous) Hall [Q(A)H], (c) quantum valley Hall (QVH), and (d) quantum (pseudo-)spin Hall (QSH) states. These phases arise from specific symmetry-breaking mechanisms, illustrating the rich topological landscape of phonons in 2D systems. Adapted from Ref. Liu et al. (2020b).
The quantum Hall family of phonons, encompassing QAH-, QSH-, and QVH- like states, can be realized by breaking specific symmetries in a honeycomb lattice Liu et al. (2020b), as illustrated in Fig. 20. The honeycomb lattice inherently supports a pair of Dirac cones at the Brillouin zone corners $K$ and $K^{\prime}$, characteristic of a 2D Dirac phonon phase, with each Dirac point associated with a quantized Berry phase of $\pi$ for a loop enclosing it. Through mechanisms such as time-reversal symmetry breaking, inversion symmetry breaking, or Kekulé distortion, the Dirac cones can be gapped, leading to the emergence of topologically distinct phases, including the QAHE, QVHE, and QSHE.
@@ III.3.1 Phononic Quantum Anomalous Hall (QAH)-like States
The quantum anomalous Hall (QAH) states in phonon systems, also known as phonon Chern insulators, represent two-dimensional gapped phases distinguished by nonzero phonon Chern numbers. These topological invariants arise from the integration of the Berry curvature, $\Omega_{z}(\mathbf{k})$, over the Brillouin zone. Crucially, in the presence of time-reversal symmetry $\mathcal{T}$, the Berry curvature must be an odd function of momentum [$\Omega_{z}(-\mathbf{k})=-\Omega_{z}(\mathbf{k})$]. This yields a zero Chern number. Consequently, the realization of a two-dimensional phonon Chern insulator inherently requires breaking $\mathcal{T}$ symmetry in the phonon system. Within the ten-fold way classification, the phonon QAH effect belongs to the “A” class, as illustrated in Figs. 3 and 4.
@@
The ordinary lattice Hamiltonian comprises a kinetic energy term $T=\sum_{i}\frac{1}{2}\dot{u}_{i}^{2}$ (where $u_{i}$ is the mass-weighted displacement of the $i$-th vibrational mode) and a harmonic potential energy term $V=\sum_{i,j}\frac{1}{2}K_{ij}u_{i}u_{j}$, both explicitly preserving time-reversal symmetry $\mathcal{T}$. To break $\mathcal{T}$ in phonon systems, a velocity-displacement bilinear term $\sum_{i,j}\eta_{ij}\dot{u}_{i}u_{j}$ was first introduced in microtubule models by Prodan and Prodan (2009a). This interaction term was later employed to study the topological phonon Hall effect Zhang et al. (2010) and topological phonon band structures Wang et al. (2015b); Huber (2016b); Süsstrunk and Huber (2016); Liu et al. (2017b).
@@
Figure 21: The honeycomb lattice model governed by $H_{0}+H_{I}+H_{T}$ (Eqs. (58), (60), (59)) hosts valley-polarized topological boundary states due to the interplay of symmetry-breaking terms $H_{I}$ (e.g., inversion-symmetry-breaking mass) and $H_{T}$ (e.g., time-reversal-symmetry-breaking flux). (a) Topological phase diagram illustrating the valley Chern number pair $(c_{K},c_{K^{\prime}})$ as function of parameters $m_{T}$ and $m_{I}$. TThe diagram is partitioned into four distinct regions by the critical lines $|m_{I}|=|m_{T}|$. Representative points A, B, C, and D in each region exhibit nearly quantized valley Chern numbers of ±$\pm\frac{1}{2}$. (b) Schematic of the 2D Brillouin zone and its projection onto the 1D boundaries. For zigzag boundary, the valleys $K$ and $K^{\prime}$ project to distinct edge $k$ points $\bar{K}$ and $\bar{K}^{\prime}$, respectively, while for armchair boundary, both valleys coalesce at the high-symmetry point $\bar{\Gamma}$. (c)-(d) Calculated spectral function of states for (c) zigzag and (d) armchair boundaries, respectively.
@@
being the Haldane-type mass term Liu et al. (2017b, 2018).
@@ III.3.2 Phononic Quantum Valley Hall (QVH)-like States
being the Semenoff-type mass term. For a small mass $m_{I}$, the Berry curvature distribution becomes predominantly concentrated around the two valleys $K$ and $K^{\prime}$, which leads to nearly quantized valley Chern numbers $c_{K}=c_{K^{\prime}}=\pm\frac{1}{2}$. The contrasting Chern numbers between the valleys facilitate the formation of a pair of valley-polarized topological boundary states. These states are characterized by a sign change in the mass $m_{I}$ across them, resulting in a shift in the valley Chern numbers $c_{K}$ and $c_{K^{\prime}}$ by ±$\pm 1$, as illustrated in Fig. 16(c). Such valley-polarized topological boundary states have been the subject of extensive research in various systems, including acoustic systems Xue et al. (2022); Zheng et al. (2020); Zhang et al. (2018c), mechanical lattices Lu et al. (2018); Zhang et al. (2018b), phononic crystals Liu et al. (2020b); Zhang et al. (2022a); Xia et al. (2022), and so on Wu et al. (2017); Liu et al. (2021b).
@@ III.3.3 Phononic Quantum Spin Hall (QSH)-like States
The quantum anomalous Hall (QAH)-like states of phonons or phonon Chern insulators requires the breaking of $\mathcal{T}$ symmetry for phonons. However, achieving $\mathcal{T}$-symmetry breaking in phonons is inherently challenging due to their charge-neutral nature and weak coupling to external electric or magnetic fields. In contrast, the quantum spin Hall (QSH) effect represents a prominent class of topological states characterized by spin-momentum-locked helical boundary states Kane and Mele (2005); Hasan and Kane (2010); Qi and Zhang (2010b, 2011). The QSH state can be interpreted as two superimposed copies of QAH states with opposite Chern numbers, ensuring cancellation of the net Chern number while preserving $\mathcal{T}$. Its effective Hamiltonian is expressed as $H_{\textrm{QSH}}=H_{0}+H_{S}$ with
@@
$s_{z}$ denotes the spin Pauli matrix. Historically, numerous pseudospin degrees of freedom have been introduced to emulate the quantum spin Hall (QSH) states within phononic systems. Examples include intricate mechanical structures, as observed in Süsstrunk and Huber (2015), and bilayer lattices, as discussed in Pal et al. (2016), among others.
The additional spin degree of freedom requires a four-band model as the minimum requirement to realize QSH. In 2015, Süsstrunk and Huber (2015) successfully constructed a mechanical topological insulator which mimics the Kane-Mele model Kane and Mele (2005) as shown in Eq. (62). Wu and Hu (2015) designed a four-band “double Dirac cone” model within a photonic honeycomb lattice characterized by a Kekulé pattern. This “double Dirac cone” arises from the folding of single Dirac cones originating from the $K$ and $K^{\prime}$ points. Liu et al. (2017a) discovered phononic pseudospins in the Kekulé lattice, exhibiting quantized pseud-angular momenta and pseudospin-polarized Berry curvature, which support pseudospin-polarized helical boundary modes that are immune to disorder-induced scattering.
@@ III.4 Phononic Stiefel-Whitney Insulator
The phononic quantum Hall (QH) family originates from the Berry curvature of phonon band structures. In 2D materials, the combined $\mathcal{PT}$ or $C_{2}\mathcal{T}$ symmetry enforces a vanishing phononic Berry curvature across the entire Brillouin zone. Nevertheless, a topological invariant—the Stiefel-Whitney (SW) number $w_{2}$ can still be defined using parity eigenvalues at time-reversal invariant momenta (TRIM) below a phononic band gap. In 3D materials, the second SW number characterize the interlocking topology of $\mathbb{Z}_{2}$ nodal lines which must exist in pairs Ahn et al. (2018); Fang et al. (2015). In contrast, the topological nodal lines with Berry flux $\pi$, i.e. a nonzero first SW number $w_{1}=1$, can exist alone Ahn et al. (2019).
@@
[…]
@@
$\hat{W}(k_{\perp})=\exp\left[-i\oint_{C}\mathrm{d}k_{\parallel}\leavevmode\nobreak\ \hat{\mathcal{A}}_{\parallel}(k_{\parallel},k_{\perp})\right]$ (64)
@@
where $\hat{\mathcal{A}}$ refers to the nonabelian Berry connection whose matrix elements is $\bm{\mathcal{A}}_{nm}=i\langle u_{n}|\bm{\nabla_{k}}u_{m}\rangle$ with $n$ and $m$ being the band indexes below the gap. $C$ is the Wilson loop parallel to $k_{\parallel}$ at fixed $k_{\perp}$. Since the Wilson operator is unitary, its eigenvalues take the form $\exp\left[-i\theta(k_{\perp})\right]$. In electronic systems, $\theta(k_{\perp})$ corresponds to the electric charge polarization, which is proportional to the Wannier centers of charge. The phase $\theta(k_{\perp})$ is only defined modulo $2\pi$, with its principal value within $[0,2\pi]$. Let the number of times $\Theta=\theta(k_{\perp})$ crosses with $\Theta=\pi$ be $N$. The parity of $N$ (even for $w_{2}=0$ and odd for $w_{2}=1$) constitutes a $\mathcal{PT}$ or $C_{2z}\mathcal{T}$-protected topological invariant, yielding the second Stiefel-Whitney number:
[…]
@@
The atomic insulator exhibits a trivial band gap with $w_{2}=0$. In contrast, a nontrivial SW insulator can be realized through a double band inversion at certain TRIM as indicated by Eq. (63). Regarding “bulk-boundary correspondence”, the SW insulator is a second-order topological insulator rather than first-order. This is because the protecting symmetry, either $\mathcal{P}$ or $C_{2z}$, is typically broken at the 1D open boundary, resulting in gapped edge states. Consequently, a 2D SW insulator with $w_{2}=1$ supports topological corner modes, characteristic of its second-order nature. As an example, Pan and Huang (2022) demonstrated the phononic SW phase in two-dimensional Xenes and their ligand-functionalized derivatives. Beyond these examples, $\mathcal{PT}$-symmetric topological phononic modes have also been explored in Kekulé lattices Liu et al. (2017a), including graphdiyne Zhang et al. (2023c), hydrogen-substituted graphdiyne Mu et al. (2022), and C3N compound Huang et al. (2023).

### v2

@@ III.3 Quantum Hall Family of Phonons
Figure 17: Various topological phonon states in 2D honeycomb lattices. (a) A honeycomb lattice hosts a 2D topological Dirac phonon phase, featuring a pair of Dirac cones at the Brillouin zone corners $K$ and $K^{\prime}$. (b)-(d) By breaking different symmetries, the Dirac cones can be gapped, leading to the emergence of topologically distinct phases, including (b) quantum (anomalous) Hall [Q(A)H], (c) quantum valley Hall (QVH), and (d) quantum (pseudo-)spin Hall (QSH) states. These phases arise from specific symmetry-breaking mechanisms, illustrating the rich topological landscape of phonons in 2D systems. Adapted from Ref. liu2020topological .
The quantum Hall family of phonons, encompassing QAH-, QSH-, and QVH- like states, can be realized by breaking specific symmetries in a honeycomb lattice liu2020topological , as illustrated in Fig. 17. The honeycomb lattice inherently supports a pair of Dirac cones at the Brillouin zone corners $K$ and $K^{\prime}$, characteristic of a 2D Dirac phonon phase, with each Dirac point associated with a quantized Berry phase of $\pi$ for a loop enclosing it. Through mechanisms such as time-reversal symmetry breaking, inversion symmetry breaking, or Kekulé distortion, the Dirac cones can be gapped, leading to the emergence of topologically distinct phases, including the QAHE, QVHE, and QSHE.
@@ III.3.1 Phononic Quantum Anomalous Hall (QAH)-like States
The quantum anomalous Hall (QAH) states in phonon systems, also known as phonon Chern insulators, represent two-dimensional gapped phases distinguished by nonzero phonon Chern numbers. These topological invariants arise from the integration of the Berry curvature, $\Omega_{z}(\mathbf{k})$, over the Brillouin zone. Crucially, in the presence of time-reversal symmetry $\mathcal{T}$, the Berry curvature must be an odd function of momentum $\Omega_{z}(-\mathbf{k})=-\Omega_{z}(\mathbf{k})$]. This yields a zero Chern number. Consequently, the realization of a two-dimensional phonon Chern insulator inherently requires breaking $\mathcal{T}$ symmetry in the phonon system. Within the ten-fold way classification, the phonon QAH effect belongs to the “A” class, as illustrated in Figs. 1 and 2.
@@
The ordinary lattice Hamiltonian comprises a kinetic energy term $T=\sum_{i}\frac{1}{2}\dot{u}_{i}^{2}$ (where $u_{i}$ is the mass-weighted displacement of the $i$-th vibrational mode) and a harmonic potential energy term $V=\sum_{i,j}\frac{1}{2}K_{ij}u_{i}u_{j}$, both explicitly preserving time-reversal symmetry $\mathcal{T}$. To break $\mathcal{T}$ in phonon systems, a Raman-type spin-lattice interaction was found to introduce the velocity-displacement coupling sheng2006theory which can lead to nonzero phonon thermal Hall conductivity Strohm2005_THE and topological phonon Hall effect Zhang2010Nov . A similar interaction term was studied in microtubles which can also lead to phonon Chern insulator phase with $C=\pm 1$ which support one-way chiral edge states of phonons prodan2009topological .
In crystalline solids, such $\mathcal{T}$-breaking effect of phonon can be realized in magnetic materials with Raman-type spin-lattice coupling sheng2006theory , Coriolis force Wang2015Jul ; wang2015topological , and molecular Berry curvature due to electron-phonon couplings saparov2022lattice ; Saito2019Dec ; Qin2011Nov which will give rise to phonon Hall effect Strohm2005_THE ; Zhang2010Nov . Recently, Refs. wang2025abinitiotheorylarge ; zhang2025electronic has successfully developed an ab initio algorithm to calculate the phonon Hall effect in a series of magnetic materials. Moreover, in the presence of such displacement-velocity coupling terms, the lattice dynamics sorely based on ordinary dynamical matrix [shown in Eq. (4)] fails to describe the phonon eigenvector and of course band topology based on it because it only characterize the properties of displacements but not velocities. This is closely related to the fact that the ordinary Newtonian or Lagrangian dynamical equation is of second order in time derivative and both initial displacements and velocities are required to determine the further state of motion. To solve this issue, Zhang2010Nov ; Susstrunk2016Aug ; Huber2016Jul ; liu2017model ; liu2018berry introduced the Hamiltonian dynamics which is of first order in time derivative.
@@
Figure 18: The honeycomb lattice model governed by $H_{0}+H_{I}+H_{T}$ (Eqs. (63), (65), (64)) hosts valley-polarized topological boundary states due to the interplay of symmetry-breaking terms $H_{I}$ (e.g., inversion-symmetry-breaking mass) and $H_{T}$ (e.g., time-reversal-symmetry-breaking flux). (a) Topological phase diagram illustrating the valley Chern number pair $(c_{K},c_{K^{\prime}})$ as function of parameters $m_{T}$ and $m_{I}$. TThe diagram is partitioned into four distinct regions by the critical lines $|m_{I}|=|m_{T}|$. Representative points A, B, C, and D in each region exhibit nearly quantized valley Chern numbers of $\pm\frac{1}{2}$. (b) Schematic of the 2D Brillouin zone and its projection onto the 1D boundaries. For zigzag boundary, the valleys $K$ and $K^{\prime}$ project to distinct edge $k$ points $\bar{K}$ and $\bar{K}^{\prime}$, respectively, while for armchair boundary, both valleys coalesce at the high-symmetry point $\bar{\Gamma}$. (c)-(d) Calculated spectral function of states for (c) zigzag and (d) armchair boundaries, respectively.
@@
being the Haldane-type mass term liu2017model ; liu2018berry .
@@ III.3.2 Phononic Quantum Valley Hall (QVH)-like States
being the Semenoff-type mass term. For a small mass $m_{I}$, the Berry curvature distribution becomes predominantly concentrated around the two valleys $K$ and $K^{\prime}$, which leads to nearly quantized valley Chern numbers $c_{K}=c_{K^{\prime}}=\pm\frac{1}{2}$. The contrasting Chern numbers between the valleys facilitate the formation of a pair of valley-polarized topological boundary states. These states are characterized by a sign change in the mass $m_{I}$ across them, resulting in a shift in the valley Chern numbers $c_{K}$ and $c_{K^{\prime}}$ by $\pm 1$, as illustrated in Fig. 13(c). Such valley-polarized topological boundary states have been the subject of extensive research in various systems, including acoustic systems Xue2022Dec ; zheng2020underwater ; zhang2018directional , mechanical lattices lu2018valley ; zhang2018topological , phononic crystals liu2020topological ; zhang2022gigahertz ; xia2022topologically , and so on wu2017direct ; liu2021valley .
@@ III.3.3 Phononic Quantum Spin Hall (QSH)-like States
The quantum anomalous Hall (QAH)-like states of phonons or phonon Chern insulators requires the breaking of $\mathcal{T}$ symmetry for phonons. However, achieving $\mathcal{T}$-symmetry breaking in phonons is inherently challenging due to their charge-neutral nature and weak coupling to external electric or magnetic fields. In contrast, the quantum spin Hall (QSH) effect represents a prominent class of topological states characterized by spin-momentum-locked helical boundary states kane2005quantum ; hasan2010colloquium ; qi2010the ; qi2011topological . The QSH state can be interpreted as two superimposed copies of QAH states with opposite Chern numbers, ensuring cancellation of the net Chern number while preserving $\mathcal{T}$. Its effective Hamiltonian is expressed as $H_{\textrm{QSH}}=H_{0}+H_{S}$ with
@@
$s_{z}$ denotes the spin Pauli matrix. Historically, numerous pseudospin degrees of freedom have been introduced to emulate the quantum spin Hall (QSH) states within phononic systems. Examples include intricate mechanical structures, as observed in susstrunk2015observation , and bilayer lattices, as discussed in pal2016helical , among others.
The additional spin degree of freedom requires a four-band model as the minimum requirement to realize QSH. In 2015, susstrunk2015observation successfully constructed a mechanical topological insulator which mimics the Kane-Mele model kane2005quantum as shown in Eq. (67). wu2015scheme designed a four-band “double Dirac cone” model within a photonic honeycomb lattice characterized by a Kekulé pattern. This “double Dirac cone” arises from the folding of single Dirac cones originating from the $K$ and $K^{\prime}$ points. liu2017pseudospins discovered phononic pseudospins in the Kekulé lattice, exhibiting quantized pseud-angular momenta and pseudospin-polarized Berry curvature, which support pseudospin-polarized helical boundary modes that are immune to disorder-induced scattering.
The 2D Dirac phonons provide a natural way of generalization to the 3D Weyl phonons. For example, by adding a $k_{z}$-dependent mass term $k_{z}\sigma_{z}$ into the effective model of 2D Dirac phonon [Eq. (63)] will lead to a 3D Weyl phonon with Chern number $C=\pm 1$, similar to the case of Eq. (39) with $N=\pm 1$. Such kind of 3D Weyl phonons with $N=\pm 1$ and $N=\pm 2$ can be realized in the chiral materials with three-fold skew rotation symmetry like Te zhang2023weyl and $\alpha$-quartz Wang2020Mar ; Lange2024Mar .
III.3.4 Transportation properties for the phonon quantum Hall states
The QAHE-like states of phonons support gapped bulk states and gapless one-way edge states simultaneously. The bulk states characterized by nonzero phonon Chern number $C$ support phonon Hall effect Zhang2010Nov . However, despite the quantization of phonon Chern number, the phonon Hall conductivity $\kappa_{xy}$ is not quantized because the phonon occupation number (Bose-Einstein statistics) is not quantized. For QSHE/QVHE like states, the gapped bulk states exhibit zero phonon Hall conductivity but nonzero phonon spin/valley Hall transport. The edge states of all the quantum Hall family all exhibit superior disorder-scattering-immune property.
Historically, the study of phonon Hall effect is initiated by experimental discovery of a magnetotransverse thermal conductivity in Tb3Ga5O12 Strohm2005_THE . Such a Hall effect is not expected to exist for phonons which do not have charge thus Lorentz force is not directly applicable. Generally, a nonzero Hall effect requires the breaking of time-reversal symmetry $\mathcal{T}$. It was soon proposed that the Raman-type spin-lattice interaction sheng2006theory which can break $\mathcal{T}$ symmetry of phonons. Later, based on the same interaction, a topological nature of phonon Hall was revealed Zhang2010Nov ; Qin2011Nov ; qin2012berry with nonzero phonon Chern number.
@@ III.4 Phononic Stiefel-Whitney Insulator
The phononic quantum Hall (QH) family originates from the Berry curvature of phonon band structures. In 2D materials, the combined $\mathcal{PT}$ or $C_{2}\mathcal{T}$ symmetry enforces a vanishing phononic Berry curvature across the entire Brillouin zone. Nevertheless, a topological invariant—the Stiefel-Whitney (SW) number $w_{2}$ can still be defined using parity eigenvalues at time-reversal invariant momenta (TRIM) below a phononic band gap. In 3D materials, the second SW number characterize the interlocking topology of $\mathbb{Z}_{2}$ nodal lines which must exist in pairs ahn2018band ; fang2015topological . In contrast, the topological nodal lines with Berry flux $\pi$, i.e. a nonzero first SW number $w_{1}=1$, can exist alone ahn2019stiefel ; bouhon2019wilson .
@@
where $\Lambda_{i}$ ($i=1,2,3,4$) denotes four TRIM in 2D Brillouin zone, and $N^{-}(\Lambda_{i})$ counts the number of negative-parity eigenmodes at $\Lambda_{i}$ below a given band gap. $\lfloor\cdot\rfloor$ refers to the floor function. The 2D gapped phononic band structures with nonzero $w_{2}$ is called phononic SW insulator or phononic real Chern insulator ding2024topological . The value of $w_{2}$ can alternatively be determined from the spectrum of the Wilson loop operator:
@@
$\hat{W}(k_{\perp})=\exp\left[-i\oint_{C}\mathrm{d}k_{\parallel}\penalty 10000\ \hat{\mathcal{A}}_{\parallel}(k_{\parallel},k_{\perp})\right]$ (69)
@@
where $\hat{\mathcal{A}}$ refers to the nonabelian Berry connection whose matrix elements is $\bm{\mathcal{A}}_{nm}=i\langle u_{n}|\bm{\nabla_{k}}u_{m}\rangle$ with $n$ and $m$ being the band indexes below the gap. $C$ is the Wilson loop parallel to $k_{\parallel}$ at fixed $k_{\perp}$. Since the Wilson operator is unitary, its eigenvalues take the form $\exp\left[-i\theta(k_{\perp})\right]$. In electronic systems, $\theta(k_{\perp})$ corresponds to the electric charge polarization, which is proportional to the Wannier centers of charge. The phase $\theta(k_{\perp})$ is only defined modulo $2\pi$, with its principal value within $0,2\pi]$. Let the number of times $\Theta=\theta(k_{\perp})$ crosses with $\Theta=\pi$ be $N$. The parity of $N$ (even for $w_{2}=0$ and odd for $w_{2}=1$) constitutes a $\mathcal{PT}$ or $C_{2z}\mathcal{T}$-protected topological invariant, yielding the second Stiefel-Whitney number:
$w_{2}=N\penalty 10000\ \mathrm{mod}\penalty 10000\ 2.$ (70)
@@
The atomic insulator exhibits a trivial band gap with $w_{2}=0$. In contrast, a nontrivial SW insulator can be realized through a double band inversion at certain TRIM as indicated by Eq. (68). Regarding “bulk-boundary correspondence”, the SW insulator is a second-order topological insulator rather than first-order. This is because the protecting symmetry, either $\mathcal{P}$ or $C_{2z}$, is typically broken at the 1D open boundary, resulting in gapped edge states. Consequently, a 2D SW insulator with $w_{2}=1$ supports topological corner modes, characteristic of its second-order nature. As an example, pan2022phononic demonstrated the phononic SW phase in two-dimensional Xenes and their ligand-functionalized derivatives. Beyond these examples, $\mathcal{PT}$-symmetric topological phononic modes have also been explored in Kekulé lattices liu2017pseudospins , including graphdiyne zhang2023pseudospin , hydrogen-substituted graphdiyne mu2022kekule , and C3N compound huang2023phononic .
The concept of Stiefel-Whitney (SW) insulator in 2D materials can be generalized to 3D by Wang2024Nov . For a generic 3D material, the band structure on two planes with $k_{z}=0$ and $k_{z}=\pi$ which preserve $\mathcal{PT}$ symmetry can be effectively described by second SW number $w_{2}(k_{z}=0)=\nu_{0}$ and $w_{2}(k_{z}=\pi)=\nu_{\pi}$. The value of $\nu_{0}$ and the relative sign between $(-1)^{\nu_{0}}$ and $(-1)^{\nu_{\pi}}$ determine the topological phases of 3D materials having “real” wavefunctions. The 3D phononic real Chern insulator (PRCI) is characterized by $(\nu_{0},\nu_{\pi})=(1,1)$, whereas $(\nu_{0},\nu_{\pi})=(1,0)$ or $(0,1)$ will indicate a phononic real semimetal (PRSM) phase. The PRSM can be characterized by a $\mathbb{Z}_{2}$ index $\nu_{\textrm{PRSM}}$ as
$(-1)^{\nu_{\textrm{PRSM}}}=(-1)^{\nu_{0}}(-1)^{\nu_{\pi}}.$ (71)
A nonzero $\nu_{\textrm{PRSM}}$ indicate a topological distinction between $k_{z}=0$ and $k_{z}=\pi$ planes so that there must exist gap closing points between $k_{z}=0$ and $k_{z}=\pi$. Such gap closing points can be either four-fold degenerate (Dirac type), triply degenerate, or two-fold degenerate nodal lines  [^1] The two-fold degenerate points cannot be Weyl points because they are prohibited by $\mathcal{PT}$ symmetry.. In the PRSM phase, the degenerate $\bm{k}$-nodes (nodal points or lines) must exist in pairs.